OpenStudy (christos):

WHERE IS MY MISTAKE?!?!? Problem: Given that sinƟ = 2/3 and Π/2 < Ɵ < Π find the exact value of cosƟ My solution: cos^2(Ɵ) + sin^2(Ɵ) = 1 sin^2(Ɵ) = 1 - cos^2(Ɵ) 4/9=1- cos^2(Ɵ) -5/9 = -cos^2(Ɵ) 5/9 = cos^2(Ɵ) cos(Ɵ) = -sqrt(5/9) Where is my mistake till now?

OpenStudy (anonymous):

for one thing, the square root of 9 is 3

OpenStudy (anonymous):

so if you were typing in an answer, it probably wanted it to look like \[\cos(\theta)=-\frac{\sqrt{5}}{3}\]

OpenStudy (anonymous):

other than that, it is all correct

OpenStudy (christos):

@satellite73 So I didnt make any mistakes at all?

geerky42 (geerky42):

No.

OpenStudy (christos):

@satellite73 confirm to me please if all of my steps above were correct, and even if they had a mistake which line?

OpenStudy (anonymous):

just draw a triangle and with 0 = 2 and h = 3.. then find a then do a/h to get cos theta

OpenStudy (anonymous):